# SimSUM – Simulated Benchmark with Structured and Unstructured Medical Records

### v3

@@ 3.2.2 Results
Table 5 compares the results obtained for all baselines. The tabular-only baselines (BN-tab and XGBoost-tab) perform consistently worse than the baselines that include text (neural-text and neural-text-tab). The evidence setting where all other features are included as evidence usually performs better than the other settings.
@@
Table 6 further breaks down the results for the neural-text classifier over the different embedding types. We clearly see that the span embeddings perform best, since this method preselects phrases that are important for the prediction of a certain symptom. While the same phrases are still implicitly embedded in the other types, the important information gets diluted as sentence embeddings are averaged, and might even be lost entirely. However, in a realistic setting, one would rarely have access to these symptom-specific span embeddings, as this would require either a manual annotation effort, or a (small) local LLM to facilitate automated annotation in a way that preserves patient privacy. We do not expect such a local LLM to achieve near the same span detection performance as GPT-4o, which we were only able to use in this case due to the artificial nature of our notes. The performance achieved with these span embeddings should therefore be seen as an upper bound for our benchmark.
@@ 3.3 Intended uses
The SimSUM dataset is primarily designed to support research on clinical information extraction in settings where tabular background variables are available. A key objective for future work is to realize the approach illustrated in Fig. 1, in which domain knowledge links structured features to textual concepts, enabling more accurate information extraction. As demonstrated in the previous section, even a simple method—concatenating tabular features with text embeddings at the input of a neural classifier—already improves symptom extraction performance, particularly on the more challenging, compact notes. This suggests that background information can help recover symptoms that are mentioned less explicitly in the text, underscoring the utility of our dataset for developing and evaluating such approaches.
Apart from this, we also foresee multiple secondary uses of the dataset. First, the dataset could facilitate research on the automation of clinical reasoning over tabular data and text, following the example of Rabaey et al. [24]. Second, it could be used to benchmark causal effect estimation methods in the presence of textual confounders, as has been done by Arno et al. [2]. This is possible thanks to the purposeful inclusion of both a treatment and outcome variable in our dataset. Third, there has been increasing interest in clinical synthetic data [7], where a set of patient characteristics is turned into a synthetic version that is meant to protect the privacy of individuals in the original dataset. Our dataset could serve as a benchmark for comparing synthetic data generation methods that jointly generate tabular variables and text [15, 5]. In short, any area of research focusing on the intersection of tabular data and text in healthcare can potentially benefit from our proposed benchmark.
@@ Acknowledgements
Appendix A Bayesian network: Additional material
@@
[…]
@@
Table 8: Train cases labeled by the expert on how many days they expect a patient to stay home (all assuming $\text{{self{\text{-}}employed}}=\text{{no}}$), with and without prescribing antibiotics. “label” indicates the expert’s estimation, while “pred” indicates the mean number of days $\lambda$ predicted by the #days models (Equation (8) and (9) in Section 2.1).
@@
Table 9: Test cases (not part of training set) labeled by the expert on how many days they expect a patient to stay home (all assuming $\text{{self{\text{-}}employed}}=\text{{no}}$), with and without prescribing antibiotics. “label” indicates the expert’s estimation, while “pred” indicates the mean number of days $\lambda$ predicted by the #days models (Equation (8) and (9) in Section 2.1).
@@
Appendix B Prompting strategy for special cases
We use an alternative prompting strategy to generate clinical notes for the patient records where none of the respiratory symptoms are present. This is the case for around one third of our dataset. In these cases, we encourage the LLM to imagine an alternative reason for the patient visit, unrelated to the respiratory domain. Fig. 5 shows a bad example of what would happen if we simply used the original prompt for these cases. Fig. 6 and Fig. 7 show the prompts we used instead, both for the cases where the patients had underlying conditions and where they did not.
@@ Appendix B Prompting strategy for special cases
Appendix C Additional example prompts
Fig. 8 and Fig. 9 show two additional example prompts.
Figure 8: Additional example prompt with generated note.
Appendix D Automated span extraction
@@
Appendix E Expert evaluation
We picked a random subset of $30$ generated notes and show them to $5$ general practitioners. All evaluators got to see the same $30$ notes, together with the prompts that were used to generate them, in a random order that differed between evaluators. All evaluators received detailed instructions on what was expected of them in the form of a PDF [^3] https://github.com/prabaey/SimSUM/blob/main/eval/Instructions_clinical_evaluation.pdf, which was orally explained by the authors. The authors then sat together with each evaluator separately to complete three example notes (different from the $30$ notes that were to be evaluated). Afterwards, the evaluators were asked to rate the notes in their own time, without the authors’ involvement. Since it is infeasible to evaluate the whole dataset of $10{,}000$ notes, we opted for a small subset of $30$ notes, each going through a relatively extensive evaluation process that considered various measures of quality (evaluators took around $5-10$ minutes to evaluate each note). We decided to show all $5$ evaluators the same set of $30$ notes, to get a broader range of expert opinions in the evaluation of each note. This also allowed for the calculation of inter-annotator agreement.
@@ Appendix E Expert evaluation
E.1 Evaluation dimensions
E.1.1 Consistency
@@ E.1.3 Clinical accuracy
E.1.4 Quality of compact version
@@
E.2 Results
E.2.1 Extended results for consistency
@@
E.2.2 Inter-Annotator Agreement
We calculate Krippendorff’s alpha with ordinal distance measure to quantify inter-annotator agreement:
@@ E.2.2 Inter-Annotator Agreement
With an alpha of 1 indicating perfect agreement, we can say that there is some agreement between the 5 evaluators. Especially for consistency, readability and realism of the physical examination, the agreement is fair. Only the ratings for content of the compact version shows no indication of agreement. We hypothesize that this is at least partially due to the high scores assigned by all evaluators (mostly 4 and 5), with a deviation from 5 being attributed to random chance rather than to a shared opinion of the evaluators. The lower standard deviations in Table 4 show that the evaluators consistently assign high scores to all aspects of the evaluation, even if they don’t agree on which particular notes deviate from these higher scores.
Appendix F Symptom predictor baselines
@@
Figure 10: Results for the neural-text baseline for the five different embedding types. We show F1-scores over the test set for both the normal and compact versions of the notes, averaging over 5 seeds. Figure best viewed in color. For numerical results, see Table 6 in Section 3.2.
F.1 BN-tab
@@
We learn a Bayesian network over the training data, providing the structure over all variables as in Fig. 2. For the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever and self-employed, we learn the conditional probability tables (CPTs) from the training data using maximum likelihood estimation (which comes down to counting co-occurrences of child and parent values for each entry in the CPT). We use the pgmpy library [1] with a K2 prior as a smoothing strategy to initialize empty CPTs.
@@
As support for learning Noisy-OR distributions is not provided in pgmpy, we learn these parameters with a custom training loop. We formulate the likelihood as in Equation (2) in Section 2.1, and learn the parameters $p_{i}$ in Equations (LABEL:eq:noisy_OR_dysp) through (LABEL:eq:noisy_OR_nasal) for the variables dyspnea, cough, pain and nasal through maximum likelihood estimation by iterating over the train set for 10 epochs, using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. To integrate the learned Noisy-OR distributions in the Bayesian network, we turn them into fully specified CPTs. To obtain these, we simply evaluate Equation (2) in Section 2.1 for all possible combinations of child and parent values. While this results in large and inefficient CPTs, the automated inference engine built into pgmpy library does not support Noisy-OR distributions directly. Note that both versions of the conditional distribution are equivalent, so we do not incur a loss in precision.
@@
Similarly, the coefficients in the logistic regression model for antibiotics and the Poisson regression model for #days are learned using maximum likelihood estimation over the training set of $8{,}000$ examples. The likelihood is expressed as in Equation (7), (8) and (9) in Section 2.1 respectively, with learnable parameters in place of each coefficient. We iterate over the train set for 15 epochs, again using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. Finally, we turn the logistic regression and Poisson regression models into CPTs by evaluating Equation (7), (8) and (9) for all combinations of parent and child values. For the variable #days, we needed to turn each discrete number of days into a category, because pgmpy only provides automated inference for Bayesian networks consisting of exclusively categorical variables. This results in a large CPT containing one row per possible number of days, which range from 0 to 15 in our training dataset, and one column for each combination of the 7 parent variables. To allow for a possible larger maximum number of days in the test set, we create a category $\geq\text{{15 days}}$, which is defined as one minus the summed probability of all other days.
@@
Once we have learned all parameters in the joint distribution, we can evaluate the Bayesian network’s ability to predict each of the symptoms. For each evidence setting (as defined in the main text), we apply variable elimination with each of the symptoms as a target variable. Looking at the causal structure in Fig. 2, we note that the model never has to marginalize over the many rows in the learned #days CPT, since it is never a target variable. This makes automated inference feasible in our case.
F.2 XGBoost-tab
@@
F.3 Neural-text
The pretrained BioLORD encoder [26] was obtained through the huggingface library. The encoder outputs 768-dimensional sentence embeddings. Since the full text did not fit into the context window, we embedded each sentence separately, and then combined them using our strategies outlined in the main text. To split the text into sentences, we used the nltk package. The settings hist, phys, mean and span all result in a text embedding of 768 dimensions, while the setting concat results in a text embedding of 2*768 dimensions.
These embeddings are fed into a linear layer with 256 neurons, followed by a ReLU activation. The hidden state is then transformed into a single output neuron, followed by a Sigmoid activation. For the classifiers that predict fever, three output neurons followed by a Softmax activation are used instead, one for each class. While the embeddings remain fixed, we learn the parameters in the hidden and output layers using cross-entropy as a loss function over the training set. We train a separate classifier for each symptom, setting and difficulty of the text (normal vs. compact). For the binary symptoms, we train for $15$ epochs using the Adam optimizer with a batch size of $100$, a learning rate of $0.001$ and weight_decay set to $51\mathrm{e}-$ 5. The classifier for $fever$ tended to collapse more easily, which is why we train it for $30$ epochs with a lower learning rate of $45\mathrm{e}-$ 4 instead. These hyperparameters were obtained using a mix of manual tuning and grid search with 5-fold cross validation over the training set.

### v4

@@ 3.3.3 Symptom predictor results
Table 4 compares the results obtained for all baselines, using the mean embedding type for the text-based classifiers. The tabular-only baselines (BN-tab and XGBoost-tab) perform consistently worse than the baselines that include text (neural-text and neural-text-tab). The evidence setting where all other features are included as evidence usually performs better than the other settings.
@@
Table 5 further breaks down the results for the neural-text classifier over the different embedding types. We clearly see that the span embeddings perform best, since this method preselects phrases that are important for the prediction of a certain symptom. While the same phrases are still implicitly embedded in the other types, the important information gets diluted as sentence embeddings are averaged, and might even be lost entirely. However, in a realistic setting, one would rarely have access to these symptom-specific span embeddings, as this would require either a manual annotation effort, or a (small) local LLM to facilitate automated annotation in a way that preserves patient privacy. We do not expect such a local LLM to achieve near the same span detection performance as GPT-4o, which we were only able to use in this case due to the artificial nature of our notes. The performance achieved with these span embeddings should therefore be seen as a best-case reference within our benchmark, rather than an achievable goal in a practical setting.
@@ 3.4 Intended uses
The primary purpose of the SimSUM dataset is to enable research on clinical information extraction in multimodal contexts, where effective integration of structured and unstructured data is required. A key objective for future work is to realize the approach illustrated in Fig. 2, in which domain knowledge links structured features to textual concepts, enabling more accurate information extraction. As demonstrated in the previous section, even a simple method—concatenating tabular features with text embeddings at the input of a neural classifier—already improves symptom extraction performance, particularly on the more challenging, compact notes. This suggests that background information can help recover symptoms that are mentioned less explicitly in the text, underscoring the utility of our dataset for developing and evaluating such approaches.
Apart from this, we also foresee multiple secondary uses of the dataset. First, the dataset could facilitate research on the automation of clinical reasoning over tabular data and text, following our previous work in rabaey2024clinical. Second, it could be used to benchmark causal effect estimation methods in the presence of textual confounders, as proposed by text_CATE and time_text_confounding. This is possible thanks to the purposeful inclusion of both a treatment and outcome variable in our dataset. Third, there has been increasing interest in clinical synthetic data [hernandez2022synthetic], where a set of patient characteristics is turned into a synthetic version that is meant to protect the privacy of individuals in the original dataset. Our dataset could serve as a benchmark for comparing synthetic data generation methods that jointly generate tabular variables and text [EHR_generation, synthetic_EMR_text, promptehr, generative_AI_synthetic]. In short, any area of research focusing on the intersection of tabular data and text in healthcare can potentially benefit from our proposed benchmark.
@@ 4 Conclusions
Ethics approval and consent to participate
Not applicable.
Consent for publication
Not applicable.
Availability of data and materials
The SimSUM dataset proposed during the current study is freely available on Github, at https://github.com/prabaey/SimSUM. This repository also contains all code necessary to reproduce the results presented in this work.
Competing interests
The authors declare that they have no competing interests.
@@ Acknowledgements
List of abbreviations
The following is a list of abbreviations that were used throughout the work.
  • EHR: Electronic health record
  • CIE: Clinical information extraction
  • BN: Bayesian network
  • DAG: Directed acyclic graph
  • CPT: Conditional probability table
  • LLM: Large language model
Appendix A Bayesian network
@@
Table 6: Test cases labeled by the expert on whether to prescribe antibiotics or not (all assume $\text{{policy}}=\text{{low}}$). “label” indicates the expert’s decision, while “pred” indicates the probability predicted by the antibiotics model, according to Equation (7).
@@
Table 7: Train cases labeled by the expert on how many days they expect a patient to stay home (all assuming $\text{{self{\text{-}}employed}}=\text{{no}}$), with and without prescribing antibiotics. “label” indicates the expert’s estimation, while “pred” indicates the mean number of days $\lambda$ predicted by the days at home models (Equation (8) and (9)).
@@
Table 8: Test cases (not part of training set) labeled by the expert on how many days they expect a patient to stay home (all assuming $\text{{self{\text{-}}employed}}=\text{{no}}$), with and without prescribing antibiotics. “label” indicates the expert’s estimation, while “pred” indicates the mean number of days $\lambda$ predicted by the days at home models (Equation (8) and (9)).
@@
Appendix B LLM prompt
B.1 Prompt details
We prompt a large language model to generate a clinical note based on the tabular variables associated with each fictional patient record. The scenario we simulate artificially is as follows. The patient goes to the primary care physician, telling them their symptoms and possible underlying conditions, along with additional context on the severity of these symptoms, when they started, and other details. The physician takes descriptive notes during this consultation, writing down the (recent) history prescribed by the patient. Then, based on the patient’s described complaints, they conduct a physical examination, writing down all findings. Both parts together then form the textual description of the patient encounter.
We now describe the different parts of the prompt.
Symptom Cause Descriptors
dyspnea asthma
attack-related, at night, in episodes, wheezing, difficulty breathing in, feeling of suffocation,
nighttime stuffiness, provoked by exercise, light, severe, not able to breathe properly, air hunger
smoking
during exercise, worse in morning, mild
COPD
chronic, worse during flare-up, worse when lying down, difficulty sleeping, air hunger
hay fever
light, mild, stuffy feeling, all closed up
pneumonia
light, mild, severe, no clear cause
cough asthma
attack-related, dry
smoking
productive, mostly in morning, during exercise, gurgling
COPD
phlegm, sputum, gurgling, worse when lying down
pneumonia
for over 7 days, light, mild, severe, non-productive at first and later purulent
common cold
prickly, irritating, dry, phlegm, sputum, light, mild, severe, constant, day and night
pain asthma
tension behind sternum
COPD
light, mild
cough
muscle pain, burning pain in trachea, burning pain in windpipe, scraping pain in trachea,
scraping pain in windpipe
pneumonia
light, mild, severe, localized on right side, localized on left side, associated with breathing
common cold
burning pain in trachea, burning pain in windpipe, scraping pain in trachea,
scraping pain in windpipe, light, mild
Model details and parameters  As a large language model, we opted for OpenAI’s GPT-4o model, using the version released in May 2024 [GPT4o]. We set the temperature to 1.2 to encourage some more variation in the notes. Before providing the case-specific prompt, we set the following system message: “You are a general practitioner, and need to summarize the patient encounter in a clinical note. Your notes are detailed and extensive.” We set the max_tokens parameter to 1000. All other parameters were set to their default values. Generating all 10,000 notes and their compact version cost around \$$150$.
B.2 Additional example prompts
Fig. 6 and Fig. 7 show two additional example prompts.
Figure 6: Additional example prompt with generated note.
B.3 Prompting strategy for special cases
Around one third of all patients in our dataset (3629 out of 10,000 patients sampled from the BN) do not experience any respiratory symptoms at all (all symptoms are “no”). If we used the same prompt as before, this would result in an unrealistic clinical note, since the note would simply list all symptoms the patient does not have, without giving an actual reason for the patient’s visit. Furthermore, there would be little variation in these notes. An example of what would happen should we use the original prompt is shown in Fig. 8.
In these cases, it makes more sense to assume that the patient visits the doctor because of a complaint unrelated to the respiratory domain, such as back pain, stomach issues, a skin rash, etc. To generate these special cases, we use a special prompt, telling the LLM the patient does not experience any of the 5 respiratory symptoms. When the patient has at least one underlying health condition (which is the case in 239 out of 3629 special cases), we add this to the prompt in the same way as before, like the example in Fig. 9. If not (i.e., for the remaining 3390 out of 3629 special cases), we tell the LLM not to mention any of those health conditions either, as shown in Fig. 10. The latter prompt asks for three clinical notes at once, encouraging the LLM to be more creative and not repeat the same scenario every time, as well as being a little more cost-effective. This is possible because of the prompt being non-specific to any of these 3390 patients. We then randomly distribute all generated notes to each tabular patient record within this subset.
@@ B.3 Prompting strategy for special cases
Appendix C Automated span extraction
@@
Appendix D Expert evaluation
We picked a random subset of $30$ generated notes and show them to $5$ general practitioners. All evaluators got to see the same $30$ notes, together with the prompts that were used to generate them, in a random order that differed between evaluators. All evaluators received detailed instructions on what was expected of them in the form of a PDF [^3] https://github.com/prabaey/SimSUM/blob/main/eval/Instructions_clinical_evaluation.pdf, which was orally explained by the authors. The authors then sat together with each evaluator separately to complete three example notes (different from the $30$ notes that were to be evaluated). Afterwards, the evaluators were asked to rate the notes in their own time, without the authors’ involvement. Since it is infeasible to evaluate the whole dataset of 10,00010{,}000 notes, we opted for a small subset of $30$ notes, each going through a relatively extensive evaluation process that considered various measures of quality (evaluators took around 5105-10 minutes to evaluate each note). We decided to show all $5$ evaluators the same set of $30$ notes, to get a broader range of expert opinions in the evaluation of each note. This also allowed for the calculation of inter-annotator agreement.
@@ Appendix D Expert evaluation
D.1 Evaluation dimensions
D.1.1 Consistency
@@ D.1.3 Clinical accuracy
D.1.4 Quality of compact version
@@
D.2 Results
D.2.1 Extended results for consistency
@@
D.2.2 Inter-Rater Reliability
Krippendorff’s alpha is a statistical measure of inter-rater reliability that quantifies the extent to which observed agreement among raters exceeds what would be expected by chance. Usually, an alpha of 0.667 is cited as the desired level in order to draw tentative conclusions [krippendorff]. We calculate Krippendorff’s alpha with ordinal distance measure between the 5 evaluators in our study, for each of the measured criteria:
@@ D.2.2 Inter-Rater Reliability
With an alpha of 1 indicating perfect agreement, and 0 indicating agreement at the level of chance, we can say that there is some agreement between the 5 evaluators. While we do not reach the desired level of 0.667, in our case we must consider that due to the high scores assigned by all evaluators (mostly 4 and 5), a deviation from 5 can more easily be attributed to random chance rather than to a shared opinion of the evaluators. The lower standard deviations in Table 2 show that the evaluators consistently assign high scores to all aspects of the evaluation, even if they do not agree on which particular notes deviate from these higher scores.
Appendix E Symptom predictor baselines
@@
Figure 11: Results for the neural-text baseline for the five different embedding types. We show F1-scores over the test set for both the normal and compact versions of the notes, averaging over 5 seeds. Figure best viewed in color. For numerical results, see Table 5 in Section $3.$
E.1 BN-tab
@@
We learn a Bayesian network over the training data, providing the structure over all variables as in Fig. 3. For the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever and self-employed, we learn the conditional probability tables (CPTs) from the training data using maximum likelihood estimation (which comes down to counting co-occurrences of child and parent values for each entry in the CPT). We use the pgmpy library [pgmpy] with a K2 prior as a smoothing strategy to initialize empty CPTs.
@@
As support for learning Noisy-OR distributions is not provided in pgmpy, we learn these parameters with a custom training loop. We formulate the likelihood as in Equation (2) in Section 2.1, and learn the parameters $p_{i}$ in Equations (LABEL:eq:noisy_OR_dysp) through (LABEL:eq:noisy_OR_nasal) for the variables dyspnea, cough, pain and nasal through maximum likelihood estimation by iterating over the train set for 10 epochs, using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. To integrate the learned Noisy-OR distributions in the Bayesian network, we turn them into fully specified CPTs. To obtain these, we simply evaluate Equation (2) in Section 2.1 for all possible combinations of child and parent values. While this results in large and inefficient CPTs, the automated inference engine built into the pgmpy library does not support Noisy-OR distributions directly. Note that both versions of the conditional distribution are equivalent, so we do not incur a loss in precision.
@@
Similarly, the coefficients in the logistic regression model for antibiotics and the Poisson regression model for days at home are learned using maximum likelihood estimation over the training set of 8,0008{,}000 examples. The likelihood is expressed as in Equation (7), (8) and (9) in Section 2.1 respectively, with learnable parameters in place of each coefficient. We iterate over the train set for 15 epochs, again using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. Finally, we turn the logistic regression and Poisson regression models into CPTs by evaluating Equation (7), (8) and (9) for all combinations of parent and child values. For the variable days at home, we needed to turn each discrete number of days into a category, because pgmpy only provides automated inference for Bayesian networks consisting of exclusively categorical variables. This results in a large CPT containing one row per possible number of days, which range from 0 to 15 in our training dataset, and one column for each combination of the 7 parent variables. To allow for a possible larger maximum number of days in the test set, we create a category $\geq\text{{15 days}}$, which is defined as one minus the summed probability of all other days.
@@
Once we have learned all parameters in the joint distribution, we can evaluate the Bayesian network’s ability to predict each of the symptoms. For each evidence setting (as defined in the main text), we apply variable elimination with each of the symptoms as a target variable. Looking at the causal structure in Fig. 3, we note that the model never has to marginalize over the many rows in the learned days at home CPT, since it is never a target variable. This makes automated inference feasible in our case.
E.2 XGBoost-tab
@@
E.3 Neural-text
The pretrained BioLORD encoder [biolord] was obtained through the huggingface library. The encoder outputs 768-dimensional sentence embeddings. Since the full text did not fit into the context window, we embedded each sentence separately, and then combined them using our strategies outlined in the main text. To split the text into sentences, we used the nltk package. The settings hist, phys, mean and span all result in a text embedding of 768 dimensions, while the setting concat results in a text embedding of 2*768 dimensions.
These embeddings are fed into a linear layer with 256 neurons, followed by a ReLU activation. The hidden state is then transformed into a single output neuron, followed by a Sigmoid activation. For the classifiers that predict fever, three output neurons followed by a Softmax activation are used instead, one for each class. While the embeddings remain fixed, we learn the parameters in the hidden and output layers using cross-entropy as a loss function over the training set. We train a separate classifier for each symptom, setting and difficulty of the text (normal vs. compact). For the binary symptoms, we train for $15$ epochs using the Adam optimizer with a batch size of $100$, a learning rate of $0.001$ and weight_decay set to $1\mathrm{e}-5$. The classifier for feverfever tended to collapse more easily, which is why we train it for $30$ epochs with a lower learning rate of $5\mathrm{e}-4$ instead. These hyperparameters were obtained using a mix of manual tuning and grid search with 5-fold cross validation over the training set.